# DiFlowDubber: Discrete Flow Matching for Automated Video Dubbing via Cross-Modal Alignment and Synchronization

### 17 Mar 2026 v2

@@ 3.4 Video Dubbing Adaptation
In this section, we describe how the pre-trained TTS framework is adapted for video dubbing, detailing the key components of the process.
@@ 3.4.1 Face-to-Prosody Mapper
The FaPro is designed to bridge facial expressions and prosodic information in speech. In video dubbing, it is crucial that facial expressions primarily determine the prosody of the dubbed speech. Using this strong correlation, we develop FaPro to generate prosodic latent representations of speech directly from facial expressions. We first process the facial features $\mathbf{v}_{\text{face}}$ using an upsampling layer, implemented as a learnable interpolation-convolution module, to expand the video sequence length to match that of the speech. Next, a ConvNeXt V2 [59] blocks are applied to enhance the temporal modeling capability of the resulting representations. Finally, a prosody predictor composed of $m$ Feed-Forward Transformer (FFT) layers iteratively extracts representations for each residual codebook of prosody tokens, where each layer models dependencies conditioned on the outputs of the previous layers. We denote the resulting prosody prior representation as: $\tilde{\mathbf{z}}_{p}=\text{FaPro}(\mathbf{v}_{\text{face}})\in\mathbb{R}^{m\times L\times D}.$
@@ 3.4.2 Synchronizer
To effectively bridge the modality gap between text, video, and speech, we propose a Synchronizer module. It enhances video-speech alignment through the dual-alignment mechanism: video-text alignment, and speech-text alignment, which together enforce robust cross-modal synchronization. In the following, we detail the key elements of this module.
@@
Leveraging the strong correlation between lip movements and phonemes, we capture cross-modal interactions using transformer blocks equipped with cross-attention, where lip features $\mathbf{v}_{\text{lip}}$ serve as queries and phoneme features $\mathbf{p}$ as keys and values. From the final transformer block, we obtain hidden states $\mathbf{h}_{\text{vt}}\in\mathbb{R}^{F\times D}$, and an attention map $\mathcal{A}_{\text{VT}}\in\mathbb{R}^{F\times N}$. To explicitly enforce alignment, we apply a contrastive loss to the attention map $\mathcal{A}_{\text{VT}}$. The durations of the phonemes are obtained from the Montreal Forced Aligner (MFA) [37], converted into frame spans via the video frame rate, and used to construct a binary alignment matrix $\mathcal{M}_{\text{VT}}\in\mathbb{R}^{F\times N}$. Each entry $m_{ij}$ indicates whether the $i$-th video frame corresponds to the $j$-th phoneme. The alignment loss is defined as:
$\scriptstyle\mathcal{L}_{\text{VT}}^{(i,j)}=-\log\frac{\sum_{i=1}^{F}\exp(m_{ij}\cdot a_{ij})/\tau}{\sum_{i=1}^{F}\exp(a_{ij})}-\log\frac{\sum_{j=1}^{N}\exp(m_{ij}\cdot a_{ij})/\tau}{\sum_{j=1}^{N}\exp(a_{ij})},$ (2)
where $a_{ij}$ denotes the attention score between the $i$-th frame and the $j$-th phoneme, and $\tau$ is temperature coefficient.
@@
We retain textual information by directly utilizing the attention weights obtained from Equation (2). Through Monotonic Alignment Search (MAS) [31], these weights are converted into phoneme-to-frame durations, allowing each phoneme embedding to be duplicated according to its assigned span. The resulting expanded sequence is then projected through a learnable transformation and added to the cross-modal hidden states $\mathbf{h}_{\text{vt}}$. We further upsample the resulting representation using a learnable interpolation-convolution module, expanding the sequence length to match the length of the speech features and projecting it into the latent speech space:
$\mathbf{h}_{s}=\text{Upsample}(\mathbf{h}_{\text{vt}}+\text{Proj}(\text{MAS}(\mathbf{p})))\in\mathbb{R}^{L\times D}.$ (3)
@@
Although the previous step aligns video and text modalities, minor misalignments can still occur after upsampling. To refine synchronization, we introduce an additional contrastive alignment between the phoneme embeddings and the speech representations. Similar to the video-text alignment, transformer blocks with cross-attention are employed, where the speech features $\mathbf{h}_{s}$ from Equation (3) serve as queries and the phoneme embeddings $\mathbf{p}$ act as keys and values. From the final transformer block, we obtain hidden states $\mathbf{h}_{\text{st}}\in\mathbb{R}^{L\times D}$ and an attention map $\mathcal{A}_{\text{ST}}\in\mathbb{R}^{L\times N}$. A binary alignment matrix $\mathcal{M}_{\text{ST}}\in\mathbb{R}^{L\times N}$ is constructed analogously to $\mathcal{M}_{\text{ST}}$ by converting the phoneme durations into speech-token spans according to the sampling rate. The phoneme-speech contrastive loss is defined as:
$\scriptstyle\mathcal{L}_{\text{ST}}^{(i,j)}=-\log\frac{\sum_{i=1}^{L}\exp(m_{ij}\cdot a_{ij})/\tau}{\sum_{i=1}^{L}\exp(a_{ij})}-\log\frac{\sum_{j=1}^{N}\exp(m_{ij}\cdot a_{ij})/\tau}{\sum_{j=1}^{N}\exp(a_{ij})},$ (4)
where $a_{ij}$ denotes the attention score between the $i$-th speech token and the $j$-th phoneme, and $\tau$ is temperature coefficient.
@@
Finally, the representation $\mathbf{h}_{\text{st}}$ is refined with a ConvNeXt V2 [59] blocks, which offers strong temporal modeling capabilities for speech-related tasks. The resulting aligned representation, denoted as $\mathbf{h}_{\text{sync}}\in\mathbb{R}^{L\times D}$, captures both the implicit alignment between video and speech, as well as the explicit alignments between video-text and speech-text pairs.
@@ 3.4.3 Content-Consistent Temporal Adaptation
CCTA facilitates the transfer of semantic content knowledge from the TTS domain and aligns it with the video to ensure temporal synchronization while preserving semantic consistency. To achieve this, we adopt the content modeling architecture and initialize its weights from the pre-trained TTS model described in Section 3.3. During adaptation, we freeze the Projection and Content Head and replace the duration prediction mechanism used in the TTS model with the Synchronizer module, as detailed in Section 3.4.2.
To ensure semantic and linguistic consistency, we impose a distillation constraint between the teacher (TTS) model and the student (video dubbing) model. The distillation loss enforces feature alignment by ensuring consistency between intermediate representations before they are passed to the content predictor. Specifically, we first fuse the global prosody representation $\tilde{\mathbf{z}}_{p}$ with the residual content information $\mathbf{h}_{\text{sync}}$. The integration is performed through a fusion layer composed of channel-wise concatenation followed by linear transformations. This simple yet effective mechanism leverages the inherent alignment between prosodic and linguistic features. The resulting fused representation is then passed transformer blocks denoted as $\mathbf{z}_{s}\in\mathbb{R}^{L\times D}$ for the student feature. Similarly, we extract the corresponding representation from the teacher model before the content predictor, denoted as $\mathbf{z}_{t}\in\mathbb{R}^{L^{\prime}\times D}$, where $L^{\prime}$ represents the temporal length of the teacher module. To align these feature spaces, we employ a cosine similarity-based distillation loss:
@@
where $\phi(\cdot)$ denotes the average pooling function, and $B$ is the batch size. Finally, a content predictor, sharing the same architecture as the prosody predictor in the FaPro module, iteratively extracts $n$ representations for each residual codebook of content tokens, denoted as $\mathbf{h}_{\text{p2c}}$, and passes them through a shared multi-layer perceptron (MLP) head to predict logits for each residual codebook, optimized using a cross-entropy (CE) loss:
@@ 3.4.4 Discrete Flow-based Prosody-Acoustic Adaptation
To effectively inherit the capability DFPA module in capturing expressive prosody and realistic acoustic characteristics learned from large-scale TTS pre-training, we initialize the denoiser with the pre-trained weights and fine-tune it with several conditioning context modifications. In particular, the global prosody prior $\tilde{\mathbf{z}}_{p}$ and the prosody-to-content correlation cues $\tilde{\mathbf{h}}_{\text{p2c}}$ are incorporated as control signals, enabling DFPA to produce diverse yet globally coherent prosody tokens together with acoustic tokens. We elaborate on the construction of the conditioning context $\mathbf{c}$ and describe how it is integrated into our framework.
@@
Given the current denoising target token sequences $\mathbf{x}_{t}$, which are obtained by concatenating the prosody token sequences $\mathbf{x}_{t}^{p}$ and the acoustic token sequences $\mathbf{x}_{t}^{a}$, we first apply dedicated embedders to encode them into latent representations: $\mathbf{e}_{t}^{p}=E_{p}(\mathbf{x}_{t}^{p})\in\mathbb{R}^{m\times L\times D},\quad\mathbf{e}_{t}^{a}=E_{a}(\mathbf{x}_{t}^{a})\in\mathbb{R}^{k\times L\times D}$. We then integrate the $\tilde{\mathbf{h}}_{\text{p2c}}$ obtained from CCTA stage by concatenation to form the complete set of attributes, including prosody, content, and acoustic representations: $\mathbf{e}_{t}=\mathbf{e}_{t}^{p}\oplus\tilde{\mathbf{h}}_{\text{p2c}}\oplus\mathbf{e}_{t}^{a}\in\mathbb{R}^{(m+n+k)\times L\times D}$, where $\oplus$ represents channel-wise concatenation. We then construct a conditioning representation $\mathbf{e}_{\text{cond}}\in\mathbb{R}^{(m+n+k)\times L\times D}$, where $\tilde{\mathbf{z}}_{p}$ serves as the conditioning signal to promote diverse and expressive prosody generation. For the remaining components, zero embeddings are concatenated to maintain consistency across all quantizer streams. Finally, this process yields the complete context representation: $\mathbf{e}=\mathbf{e}_{\text{cond}}\oplus\mathbf{e}_{t}\in\mathbb{R}^{(m+n+k)\times 2L\times D}$. This embedding is fed into the denoiser, which predicts features in $\mathbb{R}^{(m+k)\times L\times v}$ that represent the posterior distributions of the prosody and acoustic tokens, as described in Section 3.3.
@@ 3.5 Training Objectives
The total loss function $\mathcal{L}_{\text{total}}$ integrates all objectives introduced in Equations (2), (4), (6), (5), and (1). In addition, following prior works [13, 7, 9], we introduce a auxiliary Connectionist Temporal Classification (CTC) loss to encourage the output of the Synchronizer module, $\mathbf{h}_{\text{sync}}$, to retain richer linguistic information. The total loss is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{align}}+\lambda_{1}\mathcal{L}_{c}+\lambda_{2}\mathcal{L}_{\text{CTC}}+\lambda_{3}\mathcal{L}_{\text{distill}}+\lambda_{4}\mathcal{L}_{\text{DFM}},$ (7)
@@ 4.1 Experimental Setup
For zero-shot TTS pre-training, we use the LibriTTS [62]. In the video dubbing adaptation stage, our model is trained and evaluated on the Chem [25] and GRID [15] datasets. Additional details are provided in the Dataset Details section of the Supplementary Material.
@@
We employ comprehensive objective metrics to evaluate the synthesized speech. To assess temporal alignment between synthesized speech and visual content, we adopt Lip-Sync Error Confidence (LSE-C) and Lip-Sync Error Distance (LSE-D), computed using embeddings extracted from a pretrained SyncNet model [11]. For speech quality evaluation, we measure pronunciation accuracy using Word Error Rate (WER), computed with Whisper-V3 [47] as the automatic speech recognition system. Additionally, we use Speaker Encoder Cosine Similarity (SECS), following [14, 63], to quantify speaker identity consistency between the generated speech and the reference audio. Naturalness is evaluated using UTMOS [55]. To assess efficiency, we report the Real-Time Factor (RTF) under matched conditions, using a single A100 GPU and excluding video preprocessing, thereby demonstrating the computational efficiency of our discrete flow matching framework.
@@ 4.2 Main Results
As shown in Table 1, our proposed method with 128 function evaluations (NFEs) outperforms current SOTA models across most metrics on the Chem benchmark. DiFlowDubber achieves superior results in LSE-C and LSE-D, indicating stronger synchronization and demonstrating the effectiveness of our Synchronizer in enhancing temporal alignment between linguistic content and lip movements. In terms of pronunciation accuracy, our model ranks second in Setting 1.0 and achieves the best performance in Setting 2.0. Although ProDubber attains competitive WER scores in both settings, it performs poorly in lip synchronization, as reflected by its weak LSE scores, suggesting behavior closer to a TTS system rather than a true dubbing model. For speaker similarity, DiFlowDubber offers no clear advantage over the baselines, which we attribute to the shared use of identical speaker embedding model extraction in the baseline models and during SECS computation, introducing bias in favor of those systems. To further verify the true perceptual similarity, we conduct a subjective evaluation following the Mean Opinion Score (MOS) protocol under Setting 2.0. Specifically, we use MOS-Similarity (MOS-S) to evaluate speaker similarity. In this evaluation, 20 listeners rate 30 randomly selected synthesized speech samples on a 1-5 scale based on perceived similarity to the reference speech. The results show that our model achieves the highest MOS-S score among the four latest baselines, highlighting that our method achieves superior speaker faithfulness where it matters most perceptually. Regarding naturalness and speech quality, as measured by UTMOS, our method also achieves SOTA results, further validating its ability to generate natural, and expressive dubbed speech.

### 03 Apr 2026 v4

@@ 3.4 Video Dubbing Adaptation
We detail the components used to adapt the pre-trained TTS model for video dubbing.
@@ 3.4.1 Face-to-Prosody Mapper
The FaPro is designed to bridge facial expressions and prosodic information in speech. In video dubbing, it is crucial that facial expressions primarily determine the prosody of the dubbed speech. Using this strong correlation, we develop FaPro to generate prosodic latent representations of speech directly from facial expressions. We first process the facial features $\hat{\mathbf{v}}_{\text{face}}$ using an upsampling layer, implemented as a learnable interpolation-convolution module, to expand the video sequence length to match that of the speech. Next, ConvNeXt V2 [56] blocks are applied to enhance the temporal modeling capability of the resulting representations. Finally, we design a Prosody Predictor composed of $m$ Feed-Forward Transformer (FFT) layers (one per RVQ codebook of prosody) to iteratively extracts representations for each residual codebook of prosody tokens, where each layer models dependencies conditioned on the outputs of the previous layers. We denote the resulting representation as the global prosody prior: $\tilde{\mathbf{z}}_{p}=\text{FaPro}(\hat{\mathbf{v}}_{\text{face}})\in\mathbb{R}^{m\times L\times D}.$
@@ 3.4.2 Synchronizer
To effectively bridge the modality gap between text, video, and speech, we propose the Synchronizer. It enhances video-speech alignment through the dual-alignment mechanism: video-text alignment, and speech-text alignment, which together enforce robust cross-modal synchronization. In the following, we detail the key elements of this module.
@@
Leveraging the strong correlation between lip movements and phonemes, we capture cross-modal interactions using transformer blocks equipped with cross-attention, where lip features $\hat{\mathbf{v}}_{\text{lip}}$ serve as queries and phoneme features $\mathbf{p}$ as keys and values. From the final transformer block, we obtain hidden states $\mathbf{h}_{\text{vt}}\in\mathbb{R}^{F\times D}$, and an attention map $\mathcal{A}_{\text{VT}}\in\mathbb{R}^{F\times N}$. To explicitly enforce alignment, we apply a contrastive loss to the attention map $\mathcal{A}_{\text{VT}}$. The durations of the phonemes are obtained from the Montreal Forced Aligner (MFA) [35], converted into frame spans via the video frame rate, and used to construct a binary alignment matrix $\mathcal{M}_{\text{VT}}\in\mathbb{R}^{F\times N}$. Each entry $m_{ij}$ indicates whether the $i$-th video frame corresponds to the $j$-th phoneme. The alignment loss is defined as:
$\scriptstyle\mathcal{L}_{\text{VT}}=-\sum_{j=1}^{N}\log\frac{\sum_{i=1}^{F}m_{ij}\cdot\exp(a_{ij}/\tau)}{\sum_{i=1}^{F}\exp(a_{ij}/\tau)}-\sum_{i=1}^{F}\log\frac{\sum_{j=1}^{N}m_{ij}\cdot\exp(a_{ij}/\tau)}{\sum_{j=1}^{N}\exp(a_{ij}/\tau)},$ (2)
where $a_{ij}$ denotes the attention score between the $i$-th frame and the $j$-th phoneme, and $\tau$ is the temperature coefficient.
@@
We retain textual information by directly utilizing the attention weights obtained from Equation (2). Through Monotonic Alignment Search (MAS) [29], these weights are converted into phoneme-to-frame durations, allowing each phoneme embedding to be duplicated according to its assigned span, yielding $\hat{\mathbf{p}}\in\mathbb{R}^{F\times D}$. The resulting expanded sequence is then projected through a linear transformation and added to the cross-modal hidden states $\mathbf{h}_{\text{vt}}$. We further upsample the resulting representation using a learnable interpolation-convolution module, expanding the sequence length to match the length of the speech features and projecting it into the latent speech space:
$\mathbf{h}_{s}=\mathrm{Upsample}(\mathbf{h}_{\text{vt}}+\mathrm{Proj}(\hat{\mathbf{p}}))\in\mathbb{R}^{L\times D}.$ (3)
@@
Although the previous step aligns the video and text modalities, minor misalignments can still occur after upsampling. To refine synchronization, we introduce an additional contrastive alignment between the phoneme embeddings and the speech representations. Similar to the video-text alignment, transformer blocks with cross-attention are employed, where the speech features $\mathbf{h}_{s}$ from Equation (3) serve as queries and the phoneme embeddings $\mathbf{p}$ act as keys and values. From the final transformer block, we obtain hidden states $\mathbf{h}_{\text{st}}\in\mathbb{R}^{L\times D}$ and an attention map $\mathcal{A}_{\text{ST}}\in\mathbb{R}^{L\times N}$. A binary alignment matrix $\mathcal{M}_{\text{ST}}\in\mathbb{R}^{L\times N}$ is constructed analogously to $\mathcal{M}_{\text{VT}}$ by converting the phoneme durations into speech-token spans according to the sampling rate. The phoneme-speech contrastive loss is defined as:
$\scriptstyle\mathcal{L}_{\text{ST}}=-\sum_{j=1}^{N}\log\frac{\sum_{i=1}^{L}m_{ij}\cdot\exp(a_{ij}/\tau)}{\sum_{i=1}^{L}\exp(a_{ij}/\tau)}-\sum_{i=1}^{L}\log\frac{\sum_{j=1}^{N}m_{ij}\cdot\exp(a_{ij}/\tau)}{\sum_{j=1}^{N}\exp(a_{ij}/\tau)},$ (4)
where $a_{ij}$ denotes the attention score between the $i$-th speech token and the $j$-th phoneme.
@@
Finally, the representation $\mathbf{h}_{\text{st}}$ is refined with ConvNeXt V2 [56] blocks, which offer strong temporal modeling capabilities for speech-related tasks. The resulting aligned representation, denoted as $\mathbf{h}_{\text{sync}}\in\mathbb{R}^{L\times D}$, captures both the implicit alignment between video and speech, as well as the explicit alignments between video-text and speech-text pairs.
@@ 3.4.3 Content-Consistent Temporal Adaptation
CCTA facilitates the transfer of semantic content knowledge from the TTS domain and aligns it with the video to ensure temporal synchronization while preserving semantic consistency. To achieve this, we adopt the content modeling architecture and initialize its weights from the pre-trained TTS model described in Section 3.3. During adaptation, we freeze the Projection and Content Head and replace the duration prediction mechanism used in the TTS model with the Synchronizer, as detailed in Section 3.4.2.
To ensure semantic and linguistic consistency, we impose a distillation constraint between the teacher (TTS) model and the student (video dubbing) model. The distillation loss enforces feature alignment by ensuring consistency between intermediate representations before they are passed to the Content Predictor. Specifically, we first fuse the global prosody representation $\tilde{\mathbf{z}}_{p}$ with the residual content information $\mathbf{h}_{\text{sync}}$. The integration is performed through a fusion layer composed of channel-wise concatenation followed by linear transformations. This simple yet effective mechanism leverages the inherent alignment between prosodic and linguistic features. The resulting fused representation is then passed through transformer blocks denoted as $\mathbf{z}_{s}\in\mathbb{R}^{L\times D}$ for the student feature. Similarly, we extract the corresponding representation from the teacher model before the content predictor, denoted as $\mathbf{z}_{t}\in\mathbb{R}^{L^{\prime}\times D}$, where $L^{\prime}$ represents the temporal length of the teacher module. To align these feature spaces, we employ a cosine similarity-based distillation loss:
@@
where $\phi(\cdot)$ denotes the average pooling function, and $B$ is the batch size. Finally, a Content Predictor, sharing the same architecture as the Prosody Predictor in the FaPro, iteratively extracts $n$ representations for each residual codebook of content tokens, denoted as $\mathbf{h}_{\text{p2c}}\in\mathbb{R}^{n\times L\times D}$, and passes them through a Content Head, implemented as a Multi-Layer Perceptron (MLP), to predict logits for each residual codebook, optimized using a Cross-Entropy (CE) loss:
@@ 3.4.4 Discrete Flow-based Prosody-Acoustic Adaptation
To effectively inherit the capability of the DFPA module in capturing expressive prosody and realistic acoustic characteristics learned from large-scale TTS pre-training, we initialize the denoiser with the pre-trained weights and fine-tune it with several conditioning context modifications. Particularly, we replace $\mathbf{x}^{r}_{p}$ with $\tilde{\mathbf{z}}_{p}$, set $\mathbf{x}^{r}_{a}$ to zero embeddings, and substitute $\tilde{\mathbf{h}}_{c}$ with $\tilde{\mathbf{h}}_{\text{p2c}}$. These signals are incorporated as control inputs, enabling DFPA to generate diverse yet globally coherent prosody tokens alongside acoustic tokens. We further elaborate on the construction of the conditioning context $\mathbf{c}$ and describe how it is integrated into our framework.
@@
Given the current denoising target token sequences $\mathbf{x}_{t}$, which are obtained by concatenating the prosody token sequences $\mathbf{x}_{p,t}$ and the acoustic token sequences $\mathbf{x}_{a,t}$, we first apply dedicated embedders to encode them into latent representations: $\mathbf{e}_{p,t}=\mathbf{E}_{p}(\mathbf{x}_{p,t})\in\mathbb{R}^{m\times L\times D},\mathbf{e}_{a,t}=\mathbf{E}_{a}(\mathbf{x}_{a,t})\in\mathbb{R}^{k\times L\times D}$. We then integrate the $\tilde{\mathbf{h}}_{\text{p2c}}$ obtained from CCTA stage by concatenation to form the complete set of attributes, including prosody, content, and acoustic representations: $\mathbf{e}_{t}=[\mathbf{e}_{p,t};\tilde{\mathbf{h}}_{\text{p2c}};\mathbf{e}_{a,t}]\in\mathbb{R}^{(m+n+k)\times L\times D}$. We then construct a conditioning representation $\mathbf{e}_{\text{cond}}=[\tilde{\mathbf{z}}_{p};\mathbf{0};\mathbf{0}]\in\mathbb{R}^{(m+n+k)\times L\times D}$, where $\tilde{\mathbf{z}}_{p}$ serves as the conditioning signal to promote diverse and expressive prosody generation, and zero embeddings occupy the content and acoustic slots. Finally, this process yields the complete input $\mathbf{e}=[\mathbf{e}_{\text{cond}};\mathbf{e}_{t}]\in\mathbb{R}^{(m+n+k)\times 2L\times D}$. The input $\mathbf{e}$, together with a global conditioning vector $\mathbf{c}_{g}\in\mathbb{R}^{D}$ formed by the summation of the speaker embedding $\mathbf{s}$ projected onto $\mathbb{R}^{D}$ and timestep embedding $\mathbf{t}\in\mathbb{R}^{D}$, is fed into the denoiser, which predicts features in $\mathbb{R}^{(m+k)\times L\times v}$ that represent the posterior distributions of the prosody and acoustic tokens, as described in Section 3.3.
@@ 3.5 Training Objectives
The total loss of dubbing $\mathcal{L}_{\text{Dubbing}}$ integrates all objectives introduced in Equations (1), (2), (4), (5), and (6). In addition, following previous work [12, 6, 8], we introduce an auxiliary Connectionist Temporal Classification (CTC) loss to encourage the output of the Synchronizer, $\mathbf{h}_{\text{sync}}$, to retain richer linguistic information. The overall loss is defined as:
$\mathcal{L}_{\text{Dubbing}}=\lambda_{1}\mathcal{L}_{c}+\lambda_{2}\mathcal{L}_{\text{CTC}}+\lambda_{3}\mathcal{L}_{\text{distill}}+\lambda_{4}\mathcal{L}_{\text{DFM}}+\mathcal{L}_{\text{align}},$ (7)
@@ 4.1 Experimental Setup
For zero-shot TTS pre-training stage, we use the LibriTTS dataset [59]. In the video dubbing adaptation stage, our model is trained and evaluated on the Chem [24] and GRID [14] datasets. Additional details are provided in the Dataset Details section of the Supplementary Material.
@@
We employ comprehensive objective metrics to evaluate the synthesized speech. To assess temporal alignment between synthesized speech and visual content, we adopt Lip-Sync Error Confidence (LSE-C) and Lip-Sync Error Distance (LSE-D), computed using embeddings extracted from a pretrained SyncNet model [10]. For speech quality evaluation, we measure pronunciation accuracy using Word Error Rate (WER), computed with Whisper-V3 [45] as the automatic speech recognition system. Additionally, we use Speaker Encoder Cosine Similarity (SECS), following [13, 60], to quantify speaker identity consistency between the generated speech and the reference audio. Naturalness is evaluated using UTMOS [52]. To assess efficiency, we report the Real-Time Factor (RTF) under matched conditions, using a single NVIDIA A100 80GB GPU and excluding video preprocessing, thereby demonstrating the computational efficiency of our discrete flow matching framework.
@@ 4.2 Main Results
As shown in Table 1, our proposed method with 128 function evaluations (NFEs) outperforms current SOTA models across most metrics on the Chem benchmark. DiFlowDubber achieves superior results in LSE-C and LSE-D, indicating stronger synchronization and demonstrating the effectiveness of our Synchronizer in enhancing temporal alignment between linguistic content and lip movements. In terms of pronunciation accuracy, our model achieves the best performance in Setting 2.0 and ranks second in Setting 1.0. Although ProDubber attains competitive WER scores in both settings, it performs poorly in lip synchronization, as reflected by its weak LSE scores, suggesting behavior closer to a TTS system rather than a true dubbing model. For speaker similarity, DiFlowDubber offers no clear advantage over the baselines in SECS, which we attribute to the use of the same speaker embedding extractor for both baseline speech generation and SECS computation, introducing bias in favor of those systems. To further verify the true perceptual similarity, we conduct a subjective evaluation following the Mean Opinion Score (MOS) protocol under Setting 2.0. Specifically, we use MOS-Similarity (MOS-S) to evaluate speaker similarity. In this evaluation, 20 listeners rate 30 randomly selected synthesized speech samples on a 1-5 scale based on perceived similarity to the reference speech. The results show that our model achieves the highest MOS-S score among the four latest baselines, highlighting that our method achieves superior speaker faithfulness where it matters most perceptually. Regarding naturalness and speech quality, as measured by UTMOS, our method also achieves SOTA results, further validating its ability to generate natural and expressive dubbed speech.